# Whitened Score Diffusion: A Structured Prior for Imaging Inverse Problems

### v3

@@ Abstract
Conventional score-based diffusion models (DMs) may struggle with anisotropic Gaussian diffusion processes due to the required inversion of covariance matrices in the denoising score matching training objective [48]. We propose Whitened Score (WS) diffusion models, a novel framework based on stochastic differential equations that learns the Whitened Score function instead of the standard score. This approach circumvents covariance inversion, extending score-based DMs by enabling stable training of DMs on arbitrary Gaussian forward noising processes. WS DMs establish equivalence with flow matching for arbitrary Gaussian noise, allow for tailored spectral inductive biases, and provide strong Bayesian priors for imaging inverse problems with structured noise. We experiment with a variety of computational imaging tasks using the CIFAR and CelebA ($64\times 64$) datasets and demonstrate that WS diffusion priors trained on anisotropic Gaussian noising processes consistently outperform conventional diffusion priors based on isotropic Gaussian noise. Our code is open-sourced at github.com/jeffreyalido/wsdiffusion.
@@ 1 Introduction
Diffusion models (DMs) are a powerful class of generative models that implicitly learn a complex data distribution by modeling the (Stein) score function [39, 40, 16, 13, 22, 21]. The score function is then plugged into a reverse denoising process described by an ordinary differential equation (ODE) or a stochastic differential equation (SDE) to generate novel samples from noise. Typically, the forward noising process is defined by adding different levels of isotropic Gaussian noise to a clean data sample, which enables a simple and tractable denoising score matching (DSM) objective [48]. However, the DSM objective exhibits instability when the forward diffusion noise covariance is ill-conditioned or singular, as its computation requires the inversion of the covariance matrix.
Flow matching (FM) [29, 31, 1, 55] is an alternative generative modeling paradigm that reshapes an arbitrary known noise distribution into a complex data distribution according to an implicit probability path constructed by the flow. For the isotropic Gaussian case, [30, 42] established that FM and DMs are equivalent up to a rescaling of the noise parameters that define the SDE and probability paths. However, for anisotropic Gaussian noise, there exists a gap between score-based DMs and FM, where score-based DMs cannot be as easily trained for arbitrary Gaussian forward noising processes due to the necessary inversion of the covariance matrix in the conditional score [48].
A denoising DM capable of denoising structured noise is desirable in many scientific inverse problems, especially in imaging, as it may serve as a rich Bayesian prior [40, 15, 53, 24]. Imaging through fog, turbulence and scattering [3, 54, 27], wide-field microscopy [35], diffraction tomography [28, 26], optical coherence tomography (OCT) [18], interferometry [46] and many other imaging modalities have an image formation process corrupted by structured, spatially correlated noise [52, 6], in contrast to the widely assumed additive isotropic (white) Gaussian noise. Conventional DMs are trained on isotropic Gaussian noise, which may render them practically insufficient Bayesian priors for realistic use cases with correlated noise.
@@
Motivated by FM’s ability to model arbitrary probability paths and the expressiveness of diffusion priors for inverse problems, we propose Whitened Score Diffusion, a framework for learning DMs based on arbitrary Gaussian noising processes. Instead of learning the (time-dependent) score function, $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$, we learn $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, with $\mathbf{G}_{t}$ the diffusion matrix in the forward diffusion process (Fig. 1). We term our framework Whitened Score (WS) DMs, after the whitening transformation that transforms the score vector field into an isotropic vector field. This extends the current SDE framework for score-based DMs as it avoids the computation of the inverse covariance for any anisotropic Gaussian noise in DSM objective, enabling an arbitrary choice of Gaussian probability paths, similar to FM. We elaborate on the equivalence of our framework to FM and draw a connection to the reverse-time diffusion process derivation by [8], where $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ is a predictable process of the stochastic term in a reverse-time SDE.
This work presents an extension of score-based DMs to arbitrary Gaussian forward processes, bridging the gap between DMs and FM. Our framework enables a principled construction of denoising generative priors that incorporate spectral bias aligned with correlated measurement noise, leading to improved performance in inverse problems with structured noise. Empirical results on CIFAR-10 and CelebA ($64\times 64$) across several imaging tasks show consistently higher peak signal-to-noise ratio ($\mathrm{PSNR}$) reconstructions compared to conventional models trained with isotropic noise. Our contributions are: (i) a framework for training DMs that supports arbitrary Gaussian probability paths, (ii) theoretical insights and a connection to FM, and (iii) a demonstration of effective priors for imaging inverse problems under structured noise.
@@ 2.1 Score-based diffusion models
Score-based DMs are a class of generative models that estimate a probability density function $p(\mathbf{x}_{0})$ by reversing a time-dependent noising process. In continuous time, the forward noising process is described by an Itô SDE in the form of
@@
Sampling from $p(\mathbf{x}_{0})$ requires solving Eq. 2 and thereby knowing the score function, $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$. [39, 40] approximated the score function with a neural network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ by optimizing the DSM objective [48]:
@@
with $\boldsymbol{\mu}_{t}$ and $\boldsymbol{\Sigma}_{t}$ the mean and covariance of the Gaussian transition kernel $p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$. $\boldsymbol{\mu}_{t}$ and $\boldsymbol{\Sigma}_{t}$ are functions of the drift coefficient and diffusion matrix in Eq. 1 attained by solving the ODEs in Eqs. 5.50 and 5.51 in [44], creating a linearly proportional relationship as $\boldsymbol{\mu}_{t}\propto\mathbf{x}_{0}$ and $\boldsymbol{\Sigma}_{t}\propto\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$. This leads the transformed score function term, $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ in Eq. 2b to always be isotropic, as the covariance will multiply with its inverse, regardless of the diffusion coefficient, $\mathbf{G}_{t}$.
So far, all implementations of score-based DMs use uncorrelated white Gaussian noise, as the diffusion matrix $\mathbf{G}_{t}$ is diagonal [40]. Nonetheless, the current paradigm of learning the score function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ in isolation is limited to diagonal $\mathbf{G}_{t}$, as the inversion of the covariance matrix in Eq. 4 may be numerically unstable for ill-conditioned or singular $\mathbf{G}_{t}$.
Recent work by [20] shows that manipulating the spatial frequency support of the noise in the forward process shapes the inductive spectral bias of the model, thereby providing more flexible diffusion generative modeling. However, there lacks a more precise framework for training such DMs on arbitrary noising processes. Our work bridges this gap through the lens of score-based DMs in the SDE framework to create a more general picture of frequency-controlled noising processes that can be tailored to arbitrary DM tasks.
@@ 2.2 Flow matching
Flow matching (FM) [29, 31, 1] is another paradigm in generative modeling that connects a noise distribution and a data distribution with an ODE
$\frac{\mathrm{d}\phi_{t}(\mathbf{x}_{t})}{\mathrm{d}t}=\mathbf{u}_{t}(\phi_{t}(\mathbf{x}_{t}))$ (5)
@@
where $\boldsymbol{\mu}_{t}(\mathbf{x}_{0})$ and $\boldsymbol{\Sigma}_{t}(\mathbf{x}_{0})$ are the mean and covariance of the probability path $p_{t}$, and $f^{\prime}$ denotes the time derivative of $f$. Because $\boldsymbol{\Sigma}_{t}^{\prime}$ is proportional to $\boldsymbol{\Sigma}_{t}$ up to a scalar coefficient, multiplying by the inverse to yield identity [44], the functional form of $\mathbf{u}_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ allows simple and stable training of FM models with arbitrary Gaussian probability paths. This diagonal matrix-yielding multiplication currently lacks in score-based models due to the necessary inversion of the covariance matrix in Eq. 4.
We note that WS aligns with FM in the sense that both frameworks aim to enable arbitrary probability paths. In Section 3.2, we present a formal connection between WS and FM. Nevertheless, FM may require new approaches to incorporate the measurement likelihood for solving inverse problems [55], whereas our WS framework can be readily combined with existing techniques for enforcing measurement consistency (see Section 2.3 for a review).
@@ 2.3 Imaging inverse problems with diffusion model priors
Reconstructing an unknown signal $\mathbf{x}_{0}\in\mathbb{R}^{m}$ from a measurement $\mathbf{y}\in\mathbb{R}^{n}$ given a known forward model $\mathbf{y}\sim\mathcal{N}(\mathbf{A}\mathbf{x}_{0},\boldsymbol{\Sigma}_{\mathbf{y}})$—with $\boldsymbol{\Sigma}_{\mathbf{y}}\in\mathbb{R}^{n\times n}$ the covariance of the additive Gaussian noise and $\mathbf{A}\in\mathbb{R}^{m\times n}$ the measurement forward model—is a central challenge in computational imaging and scientific problems. Recent advances employ DMs as flexible priors [14], using plug-and-play schemes [56, 57, 50], likelihood-guided sampling via posterior score approximations [19, 40, 9, 11, 38, 23], Markov Chain Monte Carlo (MCMC) techniques [34, 7, 51, 53, 43, 45], variational methods [15, 33], and latent DM frameworks [36, 12, 37].
@@
The prior score is approximated by the denoising DM. However, the measurement likelihood score is intractable due to the time-dependence. Methods in [11, 5, 38, 41] make simplifying assumptions about the prior distribution, while those in [19, 10, 24, 36] treat the likelihood score approximation as an empirically designed update using the measurement as a guiding signal. All these likelihood score approximations can thus be plugged into Eq. 7 to solve the inverse problem.
A major gap in current research on imaging inverse problems is the consideration of additive structured noise. Most research on DM priors for imaging inverse problems has largely focused on scenarios with isotropic Gaussian noise, employing corresponding isotropic Gaussian denoising DMs. Recent work by [17] explored structured priors for imaging inverse problems using stochastic restoration priors achieving superior performance over conventional denoising DMs trained on isotropic Gaussian noising processes in cases involving both correlated and uncorrelated noise. However, a formal treatment of structured noise in diffusion-based frameworks lacks, which we seek to address.
@@ 3 Whitened Score Diffusion
We define our forward-time SDE with non-diagonal diffusion matrix as
@@
adopting from the variance-preserving (VP) SDE [40]. In our experiments, we constrain $\mathbf{K}$ to be in the class of circulant convolution matrices due to their ability to be implemented with the fast Fourier transform (FFT). However, our method generalizes to any $\mathbf{K}$ that is positive semidefinite. When $\mathbf{K}=\mathbf{I}$, we recover exactly the VP-SDE. The corresponding probability transition kernel of Eq. 8 is [^1] The mean and covariance of the transition kernel are solved Eqs. 5.50 and 5.51 in [44].
@@
where $\alpha_{t}=e^{-\frac{1}{2}\int_{0}^{t}\beta_{s}\mathrm{d}s}$, In general, $\alpha_{t}$ is defined as the integral of the drift coefficient from 0 to $t$, $\alpha_{t}=\int_{0}^{t}\mathbf{F}_{s}\mathrm{d}s$. By leveraging the parameterization trick for Gaussian distributions, we may rewrite Eq. 9 as the following continuous time system:

### v4

@@ Abstract
Conventional score-based diffusion models (DMs) may struggle with anisotropic Gaussian diffusion processes due to the required inversion of covariance matrices in the denoising score matching training objective [vincent_connection_2011]. We propose Whitened Score (WS) diffusion models, a novel framework based on stochastic differential equations that learns the Whitened Score function instead of the standard score. This approach circumvents covariance inversion, extending score-based DMs by enabling stable training of DMs on arbitrary Gaussian forward noising processes. WS DMs establish equivalence with flow matching for arbitrary Gaussian noise, allow for tailored spectral inductive biases, and provide strong Bayesian priors for imaging inverse problems with structured noise. We experiment with a variety of computational imaging tasks using the CIFAR, CelebA ($64\times 64$), and CelebA-HQ ($256\times 256$) datasets and demonstrate that WS diffusion priors trained on anisotropic Gaussian noising processes consistently outperform conventional diffusion priors based on isotropic Gaussian noise. Our code is open-sourced at github.com/jeffreyalido/wsdiffusion.
@@ 1 Introduction
Diffusion models (DMs) are a powerful class of generative models that implicitly learn a complex data distribution by modeling the (Stein) score function [song_generative_2020, song_score-based_2021, ho_denoising_2020, daras_soft_2022, karras_analyzing_2024, karras_elucidating_2022]. The score function is then plugged into a reverse denoising process described by an ordinary differential equation (ODE) or a stochastic differential equation (SDE) to generate novel samples from noise. Typically, the forward noising process is defined by adding different levels of isotropic Gaussian noise to a clean data sample, which enables a simple and tractable denoising score matching (DSM) objective [vincent_connection_2011]. However, the DSM objective exhibits instability when the forward diffusion noise covariance is ill-conditioned or singular, as its computation requires inverting the covariance matrix.
Flow matching (FM) [lipman_flow_2023, liu2022flowstraightfastlearning, albergo2023buildingnormalizingflowsstochastic, zhang_flow_2025] is an alternative generative modeling paradigm that reshapes an arbitrary known noise distribution into a complex data distribution according to an implicit probability path constructed by the flow. For the isotropic Gaussian case, [lipman2024flowmatchingguidecode, sun2025unifiedcontinuousgenerativemodels] established that FM and DMs are equivalent up to a rescaling of the noise parameters that define the SDE and probability paths. However, for anisotropic Gaussian noise, there exists a gap between score-based DMs and FM, where score-based DMs cannot be as easily trained for arbitrary Gaussian forward noising processes due to the necessary inversion of the covariance matrix in the conditional score [vincent_connection_2011].
A denoising DM capable of denoising structured, correlated noise is desirable in many scientific inverse problems, especially in imaging, as it may serve as a rich Bayesian prior [song_score-based_2021, Feng_2023_ICCV, zhang_improving_2024, kawar2022denoisingdiffusionrestorationmodels]. Imaging through fog, turbulence and scattering [Alido:24, zhang_imaging_2024, lin_demonstration_2025], wide-field microscopy [mockl_accurate_2020], diffraction tomography [Ling:18, Li:25], optical coherence tomography (OCT) [huang_optical_1991], interferometry [tsukui_estimating_2023] and many other imaging modalities have an image formation process corrupted by structured, spatially correlated noise [zafari_bayesian_2025, broaddus_removing_2020], in contrast to the widely assumed additive isotropic (white) Gaussian noise. Conventional DMs are trained on isotropic Gaussian noise, which may render them practically insufficient Bayesian priors for realistic use cases with correlated noise.
@@
Motivated by FM’s ability to model arbitrary probability paths and the expressiveness of diffusion priors for inverse problems, we propose Whitened Score Diffusion, a framework for learning DMs based on arbitrary Gaussian noising processes. Instead of learning the (time-dependent) score function, $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$, we learn $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, with $\mathbf{G}_{t}$ the diffusion matrix in the forward diffusion process (Fig. 1). We term our framework Whitened Score (WS) DMs, after the whitening transformation that transforms the score vector field into an isotropic vector field. This extends the current SDE framework for score-based DMs as it avoids the computation of the inverse covariance for any anisotropic Gaussian noise in DSM objective, enabling an arbitrary choice of Gaussian probability paths, similar to FM. We elaborate on the equivalence of our framework to FM and draw a connection to the reverse-time diffusion process derivation by [castanon_reverse-time_1982], where $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ is a predictable process of the stochastic term in a reverse-time SDE.
This work presents an extension of score-based DMs to arbitrary Gaussian forward processes, bridging a gap between DMs and FM. Our framework enables a principled construction of denoising generative priors that incorporate spectral bias aligned with correlated measurement noise, leading to improved performance in inverse problems with structured noise. Empirical results on CIFAR-10, CelebA ($64\times 64$), and CelebA-HQ ($256\times 256$) across several imaging tasks show consistently higher peak signal-to-noise ratio ($\mathrm{PSNR}$) reconstructions compared to conventional models trained with isotropic noise. Our contributions are: (i) a framework for training DMs that supports arbitrary Gaussian probability paths, (ii) theoretical insights and a connection to FM, and (iii) a demonstration of effective priors for imaging inverse problems under structured noise.
@@ 2.1 Score-based diffusion models
Score-based DMs are a class of generative models that estimate a probability density function $p(\mathbf{x}_{0})$ by reversing a time-dependent noising process. In continuous time, the forward noising process is described by an Itô SDE in the form of,
@@
Sampling from $p(\mathbf{x}_{0})$ requires solving Eq. 2 and thereby knowing the score function, $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$. [song_generative_2020, song_score-based_2021] approximated the score function with a neural network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ by optimizing the DSM objective [vincent_connection_2011]:
@@
with $\boldsymbol{\mu}_{t}$ and $\boldsymbol{\Sigma}_{t}$ the mean and covariance of the Gaussian transition kernel $p(\mathbf{x}_{t}\mid\mathbf{x}_{0})$. $\boldsymbol{\mu}_{t}$ and $\boldsymbol{\Sigma}_{t}$ are functions of the drift coefficient and diffusion matrix in Eq. 1 attained by solving the ODEs in Eqs. 5.50 and 5.51 in [sarkka_applied_2019], creating a linearly proportional relationship as $\boldsymbol{\mu}_{t}\propto\mathbf{x}_{0}$ and $\boldsymbol{\Sigma}_{t}\propto\mathbf{G}_{t}\mathbf{G}_{t}^{\top}$. This leads the transformed score function term, $\mathbf{G}_{t}\mathbf{G}_{t}^{\top}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ in Eq. 2b to always be isotropic, as the covariance will multiply with its inverse, regardless of the diffusion coefficient, $\mathbf{G}_{t}$.
2.2 Structured Forward Processes in Diffusion Models
Conventional score-based diffusion models (DMs) typically employ uncorrelated white Gaussian noise, corresponding to a diagonal diffusion matrix $\mathbf{G}_{t}$ [song_score-based_2021]. However, this formulation constrains the learned score function $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ to isotropic noise settings. Extending beyond diagonal $\mathbf{G}_{t}$ poses numerical challenges, as the inversion of the covariance matrix in Eq. 4 can become unstable for ill-conditioned or singular cases.
Recent studies have demonstrated that controlling the spatial frequency content of the forward noise can influence the model’s inductive spectral bias and enhance generative flexibility [jiralerspong2025shapinginductivebiasdiffusion]. Yet, a unified framework for training diffusion models under arbitrary noising processes remains lacking. Our work addresses this gap through the lens of score-based DMs in the SDE formulation, establishing a principled foundation for frequency-controlled and structured forward processes that can be adapted to diverse DM tasks.
Several prior approaches have introduced structured forward processes to improve generative expressivity. CLD [dockhorn2021score] and PSLD [pandey2023efficient] extend the state space by incorporating velocity variables, injecting noise in phase space to simplify score estimation, albeit at the cost of auxiliary dynamics. MDMs [singhal2023diffuse] and Blurring Diffusion [hoogeboom2022blurring] employ anisotropic or spatially correlated noise but require inversion of dense covariance matrices. Flexible Diffusion [harvey2022flexible] parameterizes the forward SDE to allow adaptive noise scheduling, increasing model complexity and training cost.
These advances collectively underscore the importance of moving beyond isotropic Gaussian noise while revealing practical limitations related to stability and computational overhead. Motivated by this, our proposed WS model enables arbitrary Gaussian forward processes without covariance inversion, offering a simple, stable, and general mechanism for structured generative modeling.
@@ 2.3 Flow matching
Flow matching (FM) [lipman_flow_2023, liu2022flowstraightfastlearning, albergo2023buildingnormalizingflowsstochastic] is another paradigm in generative modeling that connects a noise distribution and a data distribution with an ODE
$\frac{\mathrm{d}\phi_{t}(\mathbf{x}_{t})}{\mathrm{d}t}=\mathbf{u}_{t}(\phi_{t}(\mathbf{x}_{t})),$ (5)
@@
where $\boldsymbol{\mu}_{t}(\mathbf{x}_{0})$ and $\boldsymbol{\Sigma}_{t}(\mathbf{x}_{0})$ are the mean and covariance of the probability path $p_{t}$, and $f^{\prime}$ denotes the time derivative of $f$. Because $\boldsymbol{\Sigma}_{t}^{\prime}$ is proportional to $\boldsymbol{\Sigma}_{t}$ up to a scalar coefficient, multiplying by the inverse to yield identity [sarkka_applied_2019], the functional form of $\mathbf{u}_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ allows simple and stable training of FM models with arbitrary Gaussian probability paths. This diagonal matrix-yielding multiplication currently lacks in score-based models due to the necessary inversion of the covariance matrix in Eq. 4.
We note that WS aligns with FM in the sense that both frameworks aim to enable arbitrary probability paths. In Section 3.2, we present a formal connection between WS and FM. Nevertheless, FM may require new approaches to incorporate the measurement likelihood for solving inverse problems [zhang_flow_2025], whereas our WS framework can be readily combined with existing techniques for enforcing measurement consistency (see Section 2.4 for a review).
@@ 2.4 Imaging inverse problems with diffusion model priors
Reconstructing an unknown signal $\mathbf{x}_{0}\in\mathbb{R}^{m}$ from a measurement $\mathbf{y}\in\mathbb{R}^{n}$ given a known forward model $\mathbf{y}\sim\mathcal{N}(\mathbf{A}\mathbf{x}_{0},\boldsymbol{\Sigma}_{\mathbf{y}})$—with $\boldsymbol{\Sigma}_{\mathbf{y}}\in\mathbb{R}^{n\times n}$ the covariance of the additive Gaussian noise and $\mathbf{A}\in\mathbb{R}^{m\times n}$ the measurement forward model—is a central challenge in computational imaging and scientific problems. Recent advances employ DMs as flexible priors [daras_survey_2024], using plug-and-play schemes [zheng2025inversebenchbenchmarkingplugandplaydiffusion, zhu2023denoisingdiffusionmodelsplugandplay, wang2024dmplugpluginmethodsolving], likelihood-guided sampling via posterior score approximations [jalal_robust_2021, song_score-based_2021, choi2021ilvrconditioningmethoddenoising, chung_diffusion_2024, song2023pseudoinverseguided, kawar_snips_2021], Markov Chain Monte Carlo (MCMC) techniques [murata2023gibbsddrmpartiallycollapsedgibbs, cardoso2024monte, wu2024principledprobabilisticimagingusing, zhang_improving_2024, sun_provable_2024, trippe2023diffusion], variational methods [Feng_2023_ICCV, mardani2024a], and latent DM frameworks [rout2023solving, chung2024prompttuning, song2024solving].
@@
The prior score is approximated by the denoising DM. However, the measurement likelihood score is intractable due to the time-dependence. Methods in [chung_diffusion_2024, boys2024tweedie, song2023pseudoinverseguided, song2022solving] make simplifying assumptions about the prior distribution, while those in [jalal_robust_2021, choi_ilvr_2021, kawar2022denoisingdiffusionrestorationmodels, rout2023solving] treat the likelihood score approximation as an empirically designed update using the measurement as a guiding signal. All these likelihood score approximations can thus be plugged into Eq. 7 to solve the inverse problem.
A major gap in current research on imaging inverse problems is the consideration of additive structured noise. Most research on DM priors for imaging inverse problems has largely focused on scenarios with isotropic Gaussian noise, employing corresponding isotropic Gaussian denoising DMs. Recent work by [hu2025stochastic] explored structured priors for imaging inverse problems using stochastic restoration priors achieving superior performance over conventional denoising DMs trained on isotropic Gaussian noising processes in cases involving both correlated and uncorrelated noise. However, a formal treatment of structured noise in diffusion-based frameworks lacks, which we seek to address.
@@ 3 Whitened Score Diffusion
We define our forward-time SDE with non-diagonal diffusion matrix as,
@@
adopting from the variance-preserving (VP) SDE [song_score-based_2021]. In our experiments, we constrain $\mathbf{K}$ to be in the class of circulant convolution matrices due to their ability to be implemented with the fast Fourier transform (FFT). However, our method generalizes to any $\mathbf{K}$ that is positive semidefinite. When $\mathbf{K}=\mathbf{I}$, we recover exactly the VP-SDE. The corresponding probability transition kernel of Eq. 8 is [^1] The mean and covariance of the transition kernel are solved in Eqs. 5.50 and 5.51 in [sarkka_applied_2019].
@@
where $\alpha_{t}=e^{-\frac{1}{2}\int_{0}^{t}\beta_{s}\mathrm{d}s}$. In general, $\alpha_{t}$ is defined as the integral of the drift coefficient from 0 to $t$, $\alpha_{t}=\int_{0}^{t}\mathbf{F}_{s}\mathrm{d}s$. By leveraging the parameterization trick for Gaussian distributions, we may rewrite Eq. 9 as the following continuous time system: